The length of a rectangle is 19 m longer than the width. If the perimeter is 606 m, what are the
dimensions of the rectangle?

Found 2 solutions by mananth, ikleyn:
Answer by mananth(16946)

The length of a mrectangle is 19 m longer than the width. If the perimeter is 606 m
let width be w m
length = (w+19) m
perimeter = 2L +2w
2w +2(w+19) = 606
2w+2w+38 =606
4w+38=606
4w= 606-48
4w=558
w=139.5 m
add 19 to width to get length

Answer by ikleyn(52781)

.

            The solution by @mananth contains a  FATAL  ERROR,  which makes the answer  INCORRECT.

            I came to bring a correct solution.


let width be w m

length = (w+19) m

perimeter = 2L + 2w

2w + 2(w+19) = 606

2w + 2w + 38 = 606

4w + 38 = 606

4w = 606-38

4w = 568

w = 142 m

ANSWER.  The dimensions are 142 m and 142 + 19 = 161 m.

Solved (correctly).